### v4

@@ B.5 Sufficient Conditions for Restricted Strong Convexity
[…]
@@
(S.12$)\geq\frac{1}{2NT}\|\Theta\|_{2}^{2}\left(\min_{\alpha^{\prime}\left(\frac{x^{\prime}x}{NT}\right)\alpha=1}\mathcal{H}\left(\alpha\cdot\frac{X}{\sqrt{NT}},\mathbb{C}\right)^{2}\right).$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
00 0 $\geq\frac{1}{T\sqrt{N}}(Q_{*}(\widetilde{\beta}_{*})-Q_{*}(\beta_{0}))$
@@ B.7 Proofs for Section 4
$\widehat{\beta}^{(s+1)}-\widehat{\beta}^{\rm\,local}_{{\rm LS},R_{0}}$ $=\left[\frac{1}{NT}\,x^{\prime}\left(\mathbf{M}_{f_{0}}\otimes\mathbf{M}_{\lambda_{0}}\right)x\right]^{-1}\frac{1}{NT}\,x^{\prime}\left(\mathbf{M}_{\widehat{f}^{(s+1)}}\otimes\mathbf{M}_{\widehat{\lambda}^{(s+1)}}-\mathbf{M}_{\widehat{f}}\otimes\mathbf{M}_{\widehat{\lambda}}\right)$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\left[e-x\left(\widehat{\beta}^{(s+1)}-\beta_{0}\right)+(f_{0}\otimes\lambda_{0}){\rm vec}(\mathbf{I}_{R})\right][1+o_{P}(1)].$
@@
B.8 Proof of Section 6
Proof of Theorem 5.
@@ B.8 Proof of Section 6
[…]
@@
[…]
@@
[…]

### v5

@@ C.5 Sufficient Conditions for Restricted Strong Convexity
$\mathcal{H}\left(\alpha\cdot\frac{X}{\sqrt{NT}},\mathbb{C}\right)^{2}\geq\mu>0,\quad\text{wpa1.}$
@@
[…]
@@
$\qquad\qquad\text{s.t.}\left\|\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\right\|_{1}\leq 3\,\sqrt{2R_{0}}\,\left\|\Theta-\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\right\|_{2}.$
@@
$\qquad\qquad\qquad\qquad\text{s.t.}\qquad\left\|\Theta-\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\right\|_{2}=\omega,$
@@
$\text{s.t.}\left\|\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\right\|_{1}\leq 3\,\sqrt{2R_{0}}\,\omega.$
@@
$\leq\frac{1}{2}a_{\beta}^{2}\,T\,\sqrt{N}\,\left(\frac{c_{\rm up}}{2}+\frac{1}{T}O_{P}(1)\right)\quad{\rm wpa1}$
$\leq\frac{1}{2}a_{\beta}^{2}\,T\,\sqrt{N}\,c_{\rm up}\quad{\rm wpa1},$
@@
$0$ $\geq\frac{1}{T\sqrt{N}}(Q_{*}(\widetilde{\beta}_{*})-Q_{*}(\beta_{0}))$
@@ C.7 Proofs for Section 4
[…]
$\qquad\qquad\qquad\qquad\qquad\qquad\times\left[e-x\left(\widehat{\beta}^{(s+1)}-\beta_{0}\right)+(f_{0}\otimes\lambda_{0}){\rm vec}(\mathbf{I}_{R})\right][1+o_{P}(1)].$
@@
C.8 Proofs for Appendix B
Proof of Theorem A.1.
@@ C.8 Proofs for Appendix B
$\min_{\beta\in\mathbf{B}_{\psi}(M)}Q_{\psi}(\beta)-Q_{\psi}(\beta_{0},\Gamma_{0})>0\quad{\rm wp1},$ (S.16)
@@
$\min_{\beta\in\mathbf{B}_{\psi}(M)}Q^{*}_{\psi}(\beta)-Q^{*}_{\psi}(\beta_{0},\Gamma_{0})>0\quad{\rm wp1}.$ (S.17)
@@
$m_{it}^{*}(z)$ $:=m_{it}(z_{it}^{0})+\left(z-z_{it}^{0}\right)\,\partial_{z}m_{it}(z_{it}^{0})+b\,\int_{z_{it}^{0}}^{z}\int_{z_{it}^{0}}^{\zeta}\frac{\partial_{z^{2}}m_{it}(\xi)}{\partial_{z^{2}}\overline{m}_{it}(\xi)}d\xi d\zeta,$
where for $z<z_{it}^{0}$ the integral should be interpreted as $\int_{z_{it}^{0}}^{z}q(\zeta)d\zeta=-\int_{z}^{z_{it}^{0}}q(\zeta)d\zeta$, and analogously for the integral over $\xi$. [^27] In the quantile regression case we have $\int_{z_{it}^{0}}^{\zeta}\frac{\partial_{z^{2}}m_{it}(\xi)}{\partial_{z^{2}}\overline{m}_{it}(\xi)}d\xi=\frac{\mathbbm{1}(z_{it}^{0}<Y_{it}<\zeta)-\mathbbm{1}(\zeta<Y_{it}<z_{it}^{0})}{\partial_{z^{2}}\overline{m}_{it}(Y_{it})}.$ (S.18) Let $\overline{m}^{*}_{it}(z)=\mathbb{E}(m^{*}_{it}(z)|X)$. Our definition of $m_{it}^{*}(z)$ together with $\mathbb{E}[\partial_{z}m_{it}(z_{it}^{0})|X]=0$ imply that